### v1

@@ 3.2 Local-nonlocal coupling model
$\lambda_{2}\int_{{\Omega_{NL}}}{R_{\delta}({\mathbf{x}},{\mathbf{y}})}(u({\mathbf{x}})-u({\mathbf{y}}))\mathrm{d}{\mathbf{x}}-2\lambda_{2}\delta^{2}\int_{\Gamma}{\bar{R}_{\delta}({\mathbf{x}},{\mathbf{y}})}{\dfrac{\partial u}{\partial{\mathbf{n}}}}^{+}({\mathbf{y}})\mathrm{d}S_{\mathbf{y}}$
$=\delta^{2}\int_{\Omega_{NL}}{\bar{R}_{\delta}({\mathbf{x}},{\mathbf{y}})}f({\mathbf{y}})\mathrm{d}{\mathbf{y}}+\delta^{2}\tilde{r}_{NL,1}({\mathbf{x}}),\ {\mathbf{x}}\in\Gamma,$
@@ 5.2 Convergence analysis
$\leq$ $C\hat{B}[e_{L},e_{NL},e_{\Gamma};e_{L},e_{NL},e_{\Gamma}]+\dfrac{C}{\delta}\left\|r_{\Gamma}\right\|_{L^{2}(\Gamma)}^{2}+C\delta^{3}\left\|f\right\|_{H^{1}(\Omega)}^{2}.$
@@
Combining (35) and (42), we derive the following estimation
$\left\|e_{L}\right\|_{H^{1}({\Omega_{L}})}^{2}+\left\|e_{NL}\right\|_{L^{2}({\Omega_{NL}})}^{2}+\delta\left\|e_{\Gamma}\right\|_{L^{2}(\Gamma)}^{2}\leq C\delta^{2}\left\|f\right\|_{H^{1}(\Omega)}^{2},$
@@ 6 Discussion and Conclusion
$\int_{\Gamma_{2\delta}}\left|f({\mathbf{x}})\right|\mathrm{d}{\mathbf{x}}\leq\sum_{i=1}^{m}\int_{B({\mathbf{x}}_{i},\frac{2}{3}r_{{\mathbf{x}}_{i}})\cap{\Omega_{NL}}}\left|f({\mathbf{x}})\right|\mathrm{d}{\mathbf{x}}.$
Furthermore, for each ${\mathbf{x}}_{i}$, it is rational to assume $B({\mathbf{x}}_{i},r_{i})\cap\Gamma$ is flat and $B({\mathbf{x}}_{i},r_{i})\cap{\Omega_{NL}}$ is upper hemisphere, otherwise we can straighten the boundary. In addition, for the fixed $\delta$, we take an integer $k_{i}$ such that
$B\left({\mathbf{x}}_{i},\frac{2}{3}r_{i}\right)\subset B\left({\mathbf{x}}_{i},k_{i}\delta\right)\subset B\left({\mathbf{x}}_{i},r_{i}\right).$
When $f\in C^{1}(\overline{{\Omega_{NL}}})\cap H^{1}(\Omega)$ specially,
$\int_{B({\mathbf{x}}_{i},\frac{2}{3}r_{{\mathbf{x}}_{i}})\cap{\Omega_{NL}}}\left|f({\mathbf{x}})\right|\mathrm{d}{\mathbf{x}}$
$\leq$ $\int_{B({\mathbf{x}}_{i},k_{i}\delta)\cap\Gamma}\int_{0}^{\sqrt{k_{i}^{2}\delta^{2}-|{\mathbf{x}}^{\prime}|^{2}}}\left|f({\mathbf{x}}^{\prime},x_{n})\right|\mathrm{d}x_{n}\mathrm{d}{{\mathbf{x}}^{\prime}}$
$\leq$ $\int_{B({\mathbf{x}}_{i},k_{i}\delta)\cap\Gamma}\int_{0}^{\sqrt{k_{i}^{2}\delta^{2}-|{\mathbf{x}}^{\prime}|^{2}}}\left|f({\mathbf{x}}^{\prime},0)+\int_{0}^{x_{n}}\dfrac{\partial f}{\partial x_{n}}({\mathbf{x}}^{\prime},t)\mathrm{d}t\right|\mathrm{d}x_{n}\mathrm{d}{{\mathbf{x}}^{\prime}}$
$\leq$ $\int_{B({\mathbf{x}}_{i},k_{i}\delta)\cap\Gamma}\int_{0}^{\sqrt{k_{i}^{2}\delta^{2}-|{\mathbf{x}}^{\prime}|^{2}}}\left|f({\mathbf{x}}^{\prime},0)\right|\mathrm{d}x_{n}\mathrm{d}{\mathbf{x}}^{\prime}+\int_{0}^{x_{n}}\int_{B({\mathbf{x}}_{i},k_{i}\delta)\cap\Gamma}\int_{0}^{\sqrt{k_{i}^{2}\delta^{2}-|{\mathbf{x}}^{\prime}|^{2}}}\left|\dfrac{\partial f}{\partial x_{n}}({\mathbf{x}}^{\prime},t)\right|\mathrm{d}x_{n}\mathrm{d}{{\mathbf{x}}^{\prime}}\mathrm{d}t$
$\leq$ $C_{i}\delta\int_{B({\mathbf{x}}_{i},k_{i}\delta)\cap\Gamma}\left|f({\mathbf{x}})\right|\mathrm{d}S_{{\mathbf{x}}}+C_{i}\delta\int_{B({\mathbf{x}}_{i},k_{i}\delta)\cap{\Omega_{NL}}}\left|\nabla f({\mathbf{x}})\right|\mathrm{d}{{\mathbf{x}}}$
@@
$\int_{\Gamma_{2\delta}}\left|f({\mathbf{x}})\right|\mathrm{d}{\mathbf{x}}\leq\sum_{i=1}^{m}\int_{B({\mathbf{x}}_{i},\frac{2}{3}r_{{\mathbf{x}}_{i}})\cap{\Omega_{NL}}}\left|f({\mathbf{x}})\right|\mathrm{d}{\mathbf{x}}\leq C\delta\left\|f\right\|_{H^{1}({\Omega_{NL}})}.$
@@
$\leq$ $C\delta^{4}\left\|\nabla u\right\|_{L^{2}(\Gamma)}^{2}$
$\leq$ $C\delta^{4}\left\|u\right\|_{H^{3}({\Omega_{NL}})}^{2}.$
If we denote
$T({\mathbf{x}},{\mathbf{y}})=u({\mathbf{y}})-u({\mathbf{x}})-\sum_{i=1}^{n}\nabla^{i}u({\mathbf{x}})(y_{i}-x_{i})-\dfrac{1}{2}\sum_{i,j=1}^{n}\nabla^{i}\nabla^{j}u({\mathbf{x}})(y_{i}-x_{i})(y_{j}-x_{j}),$
We get
$\int_{{\Omega_{NL}}}{\bar{R}_{\delta}({\mathbf{x}},{\mathbf{y}})}(u({\mathbf{y}})-u({\mathbf{x}})-\nabla u({\mathbf{x}})\cdot({\mathbf{y}}-{\mathbf{x}}))\mathrm{d}{\mathbf{y}}$
$=$ $\int_{{\Omega_{NL}}}{\bar{R}_{\delta}({\mathbf{x}},{\mathbf{y}})}T({\mathbf{x}},{\mathbf{y}})\mathrm{d}{\mathbf{y}}+\dfrac{1}{2}\sum_{i,j=1}^{n}\int_{{\Omega_{NL}}}{\bar{R}_{\delta}({\mathbf{x}},{\mathbf{y}})}\nabla^{i}\nabla^{j}u({\mathbf{x}})(y_{i}-x_{i})(y_{j}-x_{j})\mathrm{d}{\mathbf{y}}.$
Because
$\int_{\Gamma}\left(\int_{{\Omega_{NL}}}{\bar{R}_{\delta}({\mathbf{x}},{\mathbf{y}})}\nabla^{i}\nabla^{j}u({\mathbf{x}})(y_{i}-x_{i})(y_{j}-x_{j})\mathrm{d}{\mathbf{y}}\right)^{2}\mathrm{d}S_{\mathbf{x}}$
$\leq$ $C\delta^{4}\int_{\Gamma}\left|\nabla^{i}\nabla^{j}u({\mathbf{x}})\right|^{2}\left(\int_{{\Omega_{NL}}}{\bar{R}_{\delta}({\mathbf{x}},{\mathbf{y}})}\mathrm{d}{\mathbf{y}}\right)^{2}\mathrm{d}S_{\mathbf{x}}$
$\leq$ $C\delta^{4}\|\nabla^{i}\nabla^{j}u({\mathbf{x}})\|^{2}_{L^{2}(\Gamma)}\leq C\delta^{4}\|u\|_{H^{3}({\Omega_{NL}})}^{2},$
all that remains to be estimated is $\int_{{\Omega_{NL}}}{\bar{R}_{\delta}({\mathbf{x}},{\mathbf{y}})}T({\mathbf{x}},{\mathbf{y}})\mathrm{d}{\mathbf{y}}$. In fact,
@@
and
$\dfrac{1}{2}\sum_{i,j=1}^{n}\nabla^{i}\nabla^{j}u({\mathbf{x}})(y_{i}-x_{i})(y_{j}-x_{j})=\sum_{i,j=1}^{n}\int_{0}^{1}\int_{0}^{1}\nabla^{i}\nabla^{j}u({\mathbf{x}})s_{1}\mathrm{d}s_{2}\mathrm{d}s_{1}(y_{i}-x_{i})(y_{j}-x_{j}).$
Thus, we have
$u({\mathbf{y}})-u({\mathbf{x}})-\sum_{i=1}^{n}\nabla^{i}u({\mathbf{x}})(y_{i}-x_{i})-\dfrac{1}{2}\sum_{i,j=1}^{n}\nabla^{i}\nabla^{j}u({\mathbf{x}})(y_{i}-x_{i})(y_{j}-x_{j})$
$=$ $\sum_{i,j=1}^{n}\int_{0}^{1}\int_{0}^{1}\int_{0}^{1}\dfrac{\mathrm{d}}{\mathrm{d}s_{3}}\nabla^{i}\nabla^{j}u({\mathbf{x}}+s_{3}s_{2}s_{1}({\mathbf{y}}-{\mathbf{x}}))s_{1}\mathrm{d}s_{3}\mathrm{d}s_{2}\mathrm{d}s_{1}(y_{i}-x_{i})(y_{j}-x_{j})$
$=$ $\sum_{i,j,k=1}^{n}\int_{0}^{1}\int_{0}^{1}\int_{0}^{1}\nabla^{i}\nabla^{j}\nabla^{k}u({\mathbf{x}}+s_{3}s_{2}s_{1}({\mathbf{y}}-{\mathbf{x}}))s_{2}s_{1}^{2}\mathrm{d}s_{3}\mathrm{d}s_{2}\mathrm{d}s_{1}\prod_{l\in\{i,j,k\}}(y_{l}-x_{l}).$
Now we get
$\int_{\Gamma}\left(\int_{\Omega_{NL}}{\bar{R}_{\delta}({\mathbf{x}},{\mathbf{y}})}T({\mathbf{x}},{\mathbf{y}})\mathrm{d}{\mathbf{y}}\right)^{2}\mathrm{d}S_{\mathbf{x}}$
$\leq$ $C\delta^{6}\int_{0}^{1}\int_{0}^{1}\int_{0}^{1}\int_{\Gamma}\int_{{\Omega_{NL}}}{\bar{R}_{\delta}({\mathbf{x}},{\mathbf{y}})}\left|\nabla^{i}\nabla^{j}\nabla^{k}u({\mathbf{x}}+s_{3}s_{2}s_{1}({\mathbf{y}}-{\mathbf{x}}))\right|^{2}\mathrm{d}{\mathbf{y}}\mathrm{d}S_{\mathbf{x}}\mathrm{d}s_{3}\mathrm{d}s_{2}\mathrm{d}s_{1}.$
For $s\in(0,1]$, take ${\mathbf{z}}={\mathbf{x}}+s({\mathbf{y}}-{\mathbf{x}})$,
$\int_{\Gamma}\int_{{\Omega_{NL}}}{\bar{R}_{\delta}({\mathbf{x}},{\mathbf{y}})}\left|\nabla^{i}\nabla^{j}\nabla^{k}u({\mathbf{x}}+s({\mathbf{y}}-{\mathbf{x}}))\right|^{2}\mathrm{d}{\mathbf{y}}\mathrm{d}S_{\mathbf{x}}$
$\leq$ $\int_{\Gamma}\int_{{\Omega_{NL}}}\alpha_{n}\delta^{-n}\bar{R}\left(\dfrac{\left|{\mathbf{x}}-{\mathbf{z}}\right|^{2}}{4s^{2}\delta^{2}}\right)\left|\nabla^{i}\nabla^{j}\nabla^{k}u({\mathbf{z}})\right|^{2}\dfrac{1}{s^{n}}\mathrm{d}{\mathbf{z}}\mathrm{d}S_{\mathbf{x}}$
$\leq$ $\int_{{\Omega_{NL}}}\left|\nabla^{i}\nabla^{j}\nabla^{k}u({\mathbf{z}})\right|^{2}\int_{\Gamma}\bar{R}_{s\delta}({\mathbf{x}},{\mathbf{z}})\mathrm{d}S_{\mathbf{x}}\mathrm{d}{\mathbf{z}}$
$\leq$ $\dfrac{C}{\delta}\left\|u\right\|_{H^{3}({\Omega_{NL}})}^{2}.$
Here we get
$\int_{\Gamma}\left(\int_{\Omega_{NL}}{\bar{R}_{\delta}({\mathbf{x}},{\mathbf{y}})}T({\mathbf{x}},{\mathbf{y}})\mathrm{d}{\mathbf{y}}\right)^{2}\mathrm{d}S_{\mathbf{x}}\leq C\delta^{5}\left\|u\right\|_{H^{3}({\Omega_{NL}})}^{2}.$
@@
When we estimate $\int_{\Omega_{NL}}{\bar{R}_{\delta}({\mathbf{x}},{\mathbf{y}})}T({\mathbf{x}},{\mathbf{y}})\mathrm{d}{\mathbf{y}}$, we have in fact assume ${\Omega_{NL}}$ is convex to help ${\mathbf{x}}+s({\mathbf{y}}-{\mathbf{x}})$ make sense. In fact, this assumption is not indispensable. We can adopt the idea of parametrization introduced in [20] to get the local convexity in the parameter space. Here we omit that complicated parametrization to make our proof highlight the essence.

### v2

@@ 3.2 Local-nonlocal coupling model
$\lambda_{2}\int_{{\Omega_{NL}}}{\bar{R}_{\delta}({\mathbf{x}},{\mathbf{y}})}(u({\mathbf{x}})-u({\mathbf{y}}))\mathrm{d}{\mathbf{x}}-2\lambda_{2}\delta^{2}\int_{\Gamma}{\bar{\bar{R}}_{\delta}({\mathbf{x}},{\mathbf{y}})}{\dfrac{\partial u}{\partial{\mathbf{n}}}}^{+}({\mathbf{y}})\mathrm{d}S_{\mathbf{y}}$
$=\delta^{2}\int_{\Omega_{NL}}{\bar{\bar{R}}_{\delta}({\mathbf{x}},{\mathbf{y}})}f({\mathbf{y}})\mathrm{d}{\mathbf{y}}+\delta^{2}\tilde{r}_{NL,1}({\mathbf{x}}),\ {\mathbf{x}}\in\Gamma,$
@@ 5.2 Convergence analysis
$\leq$ $C\hat{B}[e_{L},e_{NL},e_{\Gamma};e_{L},e_{NL},e_{\Gamma}]+\dfrac{C}{\delta}\left\|r_{\Gamma}\right\|_{L^{2}(\Gamma)}^{2}+C\delta^{3}\left\|f\right\|_{H^{1}(\Omega)}^{2}.$ (40)
@@
Combining (40)(43) and Proposition 5.1, we derive the following estimation
$\left\|e_{L}\right\|_{H^{1}({\Omega_{L}})}^{2}+\left\|e_{NL}\right\|_{H^{1}({\Omega_{NL}})}^{2}+\delta\left\|e_{\Gamma}\right\|_{L^{2}(\Gamma)}^{2}\leq C\delta^{2}\left\|f\right\|_{H^{1}(\Omega)}^{2},$
@@ 6 Discussion and Conclusion
$\int_{\Gamma_{2\delta}}\left|f({\mathbf{x}})\right|\mathrm{d}{\mathbf{x}}\leq\sum_{i=1}^{m}\int_{B({\mathbf{x}}_{i},\frac{2}{3}r_{{\mathbf{x}}_{i}})\cap\Gamma_{2\delta}}\left|f({\mathbf{x}})\right|\mathrm{d}{\mathbf{x}}.$
Furthermore, for each ${\mathbf{x}}_{i}$, it is rational to assume $B({\mathbf{x}}_{i},r_{i})\cap\Gamma$ is flat and $B({\mathbf{x}}_{i},r_{i})\cap{\Omega_{NL}}$ is upper hemisphere, otherwise we can straighten the boundary. When $f\in C^{1}(\overline{{\Omega_{NL}}})\cap H^{1}(\Omega)$ specially,
$\int_{B({\mathbf{x}}_{i},\frac{2}{3}r_{{\mathbf{x}}_{i}})\cap\Gamma_{2\delta}}\left|f({\mathbf{x}})\right|\mathrm{d}{\mathbf{x}}$
$\leq$ $\int_{B({\mathbf{x}}_{i},\frac{2}{3}r_{{\mathbf{x}}_{i}})\cap\Gamma_{2\delta}}\int_{0}^{2\delta}\left|f({\mathbf{x}}^{\prime},x_{n})\right|\mathrm{d}x_{n}\mathrm{d}{{\mathbf{x}}^{\prime}}$
$\leq$ $\int_{B({\mathbf{x}}_{i},\frac{2}{3}r_{{\mathbf{x}}_{i}})\cap\Gamma_{2\delta}}\int_{0}^{2\delta}\left|f({\mathbf{x}}^{\prime},0)+\int_{0}^{x_{n}}\dfrac{\partial f}{\partial x_{n}}({\mathbf{x}}^{\prime},t)\mathrm{d}t\right|\mathrm{d}x_{n}\mathrm{d}{{\mathbf{x}}^{\prime}}$
$\leq$ $\int_{B({\mathbf{x}}_{i},\frac{2}{3}r_{{\mathbf{x}}_{i}})\cap\Gamma_{2\delta}}\int_{0}^{2\delta}\left|f({\mathbf{x}}^{\prime},0)\right|\mathrm{d}x_{n}\mathrm{d}{\mathbf{x}}^{\prime}+\int_{0}^{x_{n}}\int_{B({\mathbf{x}}_{i},\frac{2}{3}r_{{\mathbf{x}}_{i}})\cap\Gamma_{2\delta}}\int_{0}^{2\delta}\left|\dfrac{\partial f}{\partial x_{n}}({\mathbf{x}}^{\prime},t)\right|\mathrm{d}x_{n}\mathrm{d}{{\mathbf{x}}^{\prime}}\mathrm{d}t$
$\leq$ $C_{i}\delta\int_{B({\mathbf{x}}_{i},k_{i}\delta)\cap\Gamma_{2\delta}}\left|f({\mathbf{x}})\right|\mathrm{d}S_{{\mathbf{x}}}+C_{i}\delta\int_{B({\mathbf{x}}_{i},k_{i}\delta)\cap{\Omega_{NL}}}\left|\nabla f({\mathbf{x}})\right|\mathrm{d}{{\mathbf{x}}}$
@@
$\int_{\Gamma_{2\delta}}\left|f({\mathbf{x}})\right|\mathrm{d}{\mathbf{x}}\leq\sum_{i=1}^{m}\int_{B({\mathbf{x}}_{i},\frac{2}{3}r_{{\mathbf{x}}_{i}})\cap\Gamma_{2\delta}}\left|f({\mathbf{x}})\right|\mathrm{d}{\mathbf{x}}\leq C\delta\left\|f\right\|_{H^{1}({\Omega_{NL}})}.$
@@
$\leq$ $C\delta^{4}\left\|\nabla u\right\|_{L^{2}(\Gamma)}^{2}\leq C\delta^{4}\left\|u\right\|_{H^{3}({\Omega_{NL}})}^{2}.$
For the first term,
@@
Noticing that, with $\mathbf{z}={\mathbf{x}}+s({\mathbf{y}}-{\mathbf{x}})$,
$\delta^{4}\int_{\Gamma}\int_{{\Omega_{NL}}}{\bar{R}_{\delta}({\mathbf{x}},{\mathbf{y}})}\left|\nabla^{i}\nabla^{j}u({\mathbf{x}}+s({\mathbf{y}}-{\mathbf{x}}))\right|^{2}\mathrm{d}{\mathbf{y}}\mathrm{d}S_{\mathbf{x}}$
$\leq$ $\delta^{4}\int_{\Gamma}\int_{{\Omega_{NL}}}\delta^{-n}\bar{R}\left(\frac{|{\mathbf{x}}-\mathbf{z}|^{2}}{4s^{2}\delta^{2}}\right)\left|\nabla^{i}\nabla^{j}u(\mathbf{z})\right|^{2}\frac{1}{s^{n}}\mathrm{d}\mathbf{z}\mathrm{d}S_{\mathbf{x}}$
$=$ $\delta^{4}\int_{\Gamma}\int_{{\Omega_{NL}}}\bar{R}_{s\delta}({\mathbf{x}},\mathbf{z})\left|\nabla^{i}\nabla^{j}u(\mathbf{z})\right|^{2}\mathrm{d}\mathbf{z}\mathrm{d}S_{\mathbf{x}}$
$=$ $\delta^{4}\int_{\Gamma}\int_{\Gamma_{2s\delta}}\bar{R}_{s\delta}({\mathbf{x}},\mathbf{z})\left|\nabla^{i}\nabla^{j}u(\mathbf{z})\right|^{2}\mathrm{d}\mathbf{z}\mathrm{d}S_{\mathbf{x}}$
$\leq$ $\delta^{4}\frac{1}{s\delta}\int_{\Gamma_{2s\delta}}\left|\nabla^{i}\nabla^{j}u(\mathbf{z})\right|^{2}\mathrm{d}\mathbf{z}$
$\leq$ $\delta^{4}\left\|u\right\|_{H^{3}({\Omega_{NL}})}.$
In the last inequality, we used an estimation
$\left\|u\right\|_{L^{2}(\Gamma_{\sigma})}\leq C\sigma\left\|u\right\|_{H^{1}({\Omega_{NL}})},\quad\text{for }u\in H^{1}({\Omega_{NL}}),\sigma<<1,$
which can be proved with a same method in Appendix A. Now we can get
$\left\|\int_{{\Omega_{NL}}}{\bar{R}_{\delta}({\mathbf{x}},{\mathbf{y}})}(u({\mathbf{x}})-u({\mathbf{y}})-\nabla u({\mathbf{x}})\cdot({\mathbf{y}}-{\mathbf{x}}))\mathrm{d}{\mathbf{y}}\right\|_{L^{2}({\Omega_{NL}})}\leq C\delta^{2}\left\|u\right\|_{H^{3}({\Omega_{NL}})}.$
@@
When we estimate the first term, we have in fact assume ${\Omega_{NL}}$ is convex to help ${\mathbf{x}}+s({\mathbf{y}}-{\mathbf{x}})$ make sense. In fact, this assumption is not indispensable. We can adopt the idea of parametrization introduced in [20] to get the local convexity in the parameter space. Here we omit that complicated parametrization to make our proof highlight the essence.